# Twisted Fourier transforms on non-Kac compact quantum groups

### v1

@@ Abstract.
We introduce an analytic family of twisted Fourier transforms $\left\{\mathcal{F}^{(x)}_{p}\right\}_{x\in\mathbb{R},p\in[1,2)}$ for non-Kac compact quantum groups and establish a sharpened form of the Hausdorff-Young inequality in the range $0\leq x\leq 1$. As an application, we derive a stronger form of the twisted rapid decay property for polynomially growing non-Kac discrete quantum groups, including the duals of the Drinfeld-Jimbo $q$-deformations. Furthermore, we prove that the range $0\leq x\leq 1$ is both necessary and sufficient for the boundedness of $\mathcal{F}^{(x)}_{p}$ under the assumption of sub-exponential growth on the dual discrete quantum group. We also show that the range of boundedness of $\mathcal{F}^{(x)}_{p}$ can be strictly extended beyond $[0,1]$ for certain non-Kac and non-coamenable free orthogonal quantum groups.
@@ 1. Introduction
where the modular operator $Q=(Q_{\alpha})_{\alpha\in\text{Irr}(\mathbb{G})}$ is defined in Subsection 2.1. A central question addressed in this paper is whether the twisted Fourier transform $\mathcal{F}^{(x)}_{p}$ extends to a bounded map from $L^{p}(\mathbb{G})$ to $\ell^{p^{\prime}}(\widehat{\mathbb{G}})$ where $\widehat{\mathbb{G}}$ denotes the dual discrete quantum group of $\mathbb{G}$. The main result of this section is the establishment of the following strong Hausdorff-Young inequality (Theorem 3.6) for all $21\leq p<$ 2:
@@
In contrast, for $p>2$, we show that the twisted Fourier transform $\mathcal{F}^{(x)}_{p}$ extends to a bounded map from $L^{p}(\mathbb{G})$ to $\ell^{p^{\prime}}(\widehat{\mathbb{G}})$ only when $\mathbb{G}$ is finite-dimensional (Proposition 3.3), similarly to the classical situation.
@@
$I(\mathbb{G},p)$ =$=\left\{x\in\mathbb{R}~{}\Bigr{|}~{}\mathcal{F}^{(x)}_{p}:L^{p}(\mathbb{G})\rightarrow\ell^{p^{\prime}}(\widehat{\mathbb{G}})\text{ is bounded}\right\}.$ (1.4)
Note that Theorem 3.6 guarantees that $[0,1]\subseteq I(\mathbb{G},p)$ for any compact quantum group $\mathbb{G}$ and $21\leq p<$ 2. Subsection 5.1 focuses on non-Kac and coamenable compact quantum groups, and shows that equality $[0,1]=I(\mathbb{G},p)$ holds for all $21\leq p<$ 2 under the assumption of sub-exponential growth on $\widehat{\mathbb{G}}$ (Theorem 5.2). This class includes the Drinfeld-Jimbo $q$-deformations $G_{q}$. Subsection 5.2 focuses on non-Kac and non-coamenable free orthogonal quantum groups $O_{F}^{+}$. Using the recently established Haagerup inequality on $O_{F}^{+}$ [BVY21, You22], we show that the inclusion $[0,1]\subsetneq I(O_{F}^{+},p)$ is strict for certain non-Kac and non-coamenable free orthogonal quantum groups $O_{F}^{+}$ satisfying a technical assumption $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$ where $d_{1}=\text{Tr}(F^{*}F)$ (Corollary 5.7). A summary of the main conclusions regarding $I(\mathbb{G},p)$ is presented in the following table:
@@
$21\leq p<$ 2 $I(\mathbb{G},p)=\mathbb{R}$ $I(G_{q},p)=[0,1]$ $I(O_{F}^{+},p)\supsetneq[0,1]$
@@
$2<p\leq\infty$ $I(\mathbb{G},p)=\left\{\begin{array}[]{ll}\mathbb{R}&,~{}\mathbb{G}\text{ is finite-dimensional}\\ \emptyset&,~{}\mathbb{G}\text{ is infinite-dimensional}\end{array}\right.$ (Proposition 3.3)
@@ 2.1. Compact quantum groups
[KS18, Section 6] Let $\mathbb{G}$ be a compact matrix quantum group of non-Kac type and let ||$\cdot$|| ⋅ | be the natural length function on $\text{Irr}(\mathbb{G})$. Then there exists a sequence $(\alpha_{k})_{k\in\mathbb{N}}\subseteq\text{Irr}(\mathbb{G})$ such that
@@ 2.2. Noncommutative $L^{p}$-spaces
$[\iota(a)](x)=h(xa),~{}x\in L^{\infty}(\mathbb{G}).$ (2.14)
@@
Let us denote by $S=\left\{z\in\mathbb{C}:0<\text{Re}(z)<1\right\}$ and by $\mathcal{C}(A)$ the set of all continuous bounded functions $f:\overline{S}\rightarrow M_{n}(\mathbb{C})$ such that $f|_{S}:S\rightarrow M_{n}$ is analytic and $f(\frac{1}{p})=A$. The finite dimensional Schatten $p$-space $S^{p}_{n}$ is the complex interpolation space $(M_{n},S^{1}_{n})_{\theta}$ with $\theta=\frac{1}{p}$ in the sense that
@@ 2.3. Discrete quantum group and Fourier transform
For any $A=(A(\alpha))_{\alpha\in\text{Irr}(\mathbb{G})}\in\ell^{\infty}(\widehat{\mathbb{G}})$, we denote the support of $A$ by
@@
for all $A=(A(\alpha))_{\alpha\in\text{Irr}(\mathbb{G})}\in\ell^{\infty}(\widehat{\mathbb{G}})_{+}$.
@@
and the $\ell^{p}$-norm of $A=(A(\alpha))_{\alpha\in\text{Irr}(\mathbb{G})}\in\ell^{p}(\widehat{\mathbb{G}})$ is explicitly given by
@@
Similarly as in Subsection 2.2, a natural dual pairing between $\ell^{p}(\widehat{\mathbb{G}})$ and $\ell^{p^{\prime}}(\widehat{\mathbb{G}})$ is given by
@@ 3. Twisted Fourier transforms
For any $x\in\mathbb{R}$ and $p\in[1,\infty]$, let us consider a linear map $\mathcal{F}^{(x)}_{p}:\text{Pol}(\mathbb{G})\rightarrow c_{00}(\widehat{\mathbb{G}})$ given by
@@
$\mathcal{F}^{(x)}_{p}:L^{p}(\mathbb{G})\rightarrow\ell^{p^{\prime}}(\widehat{\mathbb{G}}).$ (3.2)
@@
$I(\mathbb{G},p)=\left\{x\in\mathbb{R}:\mathcal{F}^{(x)}_{p}:L^{p}(\mathbb{G})\rightarrow\ell^{p^{\prime}}(\widehat{\mathbb{G}})\text{ is bounded}\right\}.$ (3.3)
@@ 3.1. Unboundedness for the cases $p>2$
$I(\mathbb{G},p)=\left\{\begin{array}[]{ll}\mathbb{R}&,~{}\text{if }\mathbb{G}\text{ is finite-dimensional}\\ \emptyset&,~{}\text{if }\mathbb{G}\text{ is infinite-dimensional}\end{array}\right..$ (3.18)
@@
(1) Let us suppose that $\mathbb{G}$ is of Kac type and $p\geq 2$. Then the Fourier transform $f\in\text{Pol}(\mathbb{G})\mapsto\widehat{f}\in c_{00}(\widehat{\mathbb{G}})$ extends to a bounded linear map from $L^{p}(\mathbb{G})$ into $\ell^{p^{\prime}}(\widehat{\mathbb{G}})$ if and only if $\mathbb{G}$ is finite-dimensional by [You18, Subsection 3.5] whose arguments rely essentially on the theory of random Fourier series.
@@
In this case, the numerator is given by $[(Q_{\alpha})_{jj}(Q_{\alpha})_{ii}^{-1}]^{(\frac{1}{2}-\frac{1}{p})x}\cdot\left(\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right)^{\frac{1}{p}}$ by Lemma 3.2, and we have
@@
3.2. Boundedness for the cases $21\leq p<$ 2 and $0\leq x\leq 1$
In this Subsection, we will focus on the cases $21\leq p<$ 2. If $\mathbb{G}$ is of Kac type, then we have $\mathcal{F}^{(x)}_{p}=\mathcal{F}^{(0)}_{p}$ for all $x\in\mathbb{R}$, so we obtain $\mathbb{R}=I(\mathbb{G},p)$ for all $21\leq p<$ 2 thanks to the standard Hausdorff-Young inequality.
@@ 3.2. Boundedness for the cases $21\leq p<$ 2 and $0\leq x\leq 1$
The main result of this Section is that $[0,1]\subseteq I(\mathbb{G},p)$ for all $21\leq p<$ 2. Let us begin with the case $p=1$ in the following theorem.
@@
In particular, we have $[0,1]\subseteq I(\mathbb{G},1)$.
@@
As the second step, let us interpolate the two extremal cases $x=0$ and $x=1$ as follows. Let us fix $\alpha\in\text{Irr}(\mathbb{G})$ and consider linear operators $T_{z}:\text{Pol}(\mathbb{G})\rightarrow\ell^{\infty}(\widehat{\mathbb{G}})$ given by
@@
For the given real number $x\in I(\mathbb{G},1)$, let us consider a linear map $\Phi_{x}:\text{Pol}(\mathbb{G})\rightarrow c_{00}(\widehat{\mathbb{G}})$ given by
@@
Let $21\leq p<$ 2. For each $\alpha\in\text{Irr}(\mathbb{G})$, let us take indices $i=i(\alpha),j=j(\alpha)$ such that $(Q_{\alpha})_{ii}\geq(Q_{\alpha})_{jj}$. Then, for any $f\sim\sum_{\alpha\in\text{Irr}(\mathbb{G})}a_{\alpha}u^{\alpha}_{ij}\in L^{p}(\mathbb{G})$ and $x\leq y$, we have
@@
=$=\left\{\begin{array}[]{lll}\left(\sum_{\alpha\in\text{Irr}(\mathbb{G})}|a_{\alpha}|^{p^{\prime}}\left[\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right]^{p^{\prime}-1}\left[(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}\right]^{\frac{p^{\prime}}{2}-1}\right)^{\frac{1}{p^{\prime}}}&,~{}1<p<2\\ \sup_{\alpha\in\text{Irr}(\mathbb{G})}\frac{|a_{\alpha}|}{d_{\alpha}}[(Q_{\alpha})_{jj}(Q_{\alpha})_{ii}]^{-\frac{1}{2}}&,~{}p=1.\end{array}\right.$ (3.54)
@@
=$=\left\{\begin{array}[]{ll}\left(\sum_{\alpha\in\text{Irr}(\mathbb{G})}\left[(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}\right]^{(\frac{p^{\prime}}{2}-1)x}\left[\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right]^{p^{\prime}-1}\left|a_{\alpha}\right|^{p^{\prime}}\right)^{\frac{1}{p^{\prime}}}&,~{}1<p<2\\ \sup_{\alpha\in\text{Irr}(\mathbb{G})}\left[(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}\right]^{\frac{x}{2}}\cdot\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}|a_{\alpha}|&p=1.\end{array}\right.$ (3.58)
@@
Let $\mathbb{G}=SU_{q}(2)$ with $10<q<$ 1. Then $\text{Irr}(SU_{q}(2))$ is identified with $\left\{u^{(k)}\right\}_{k\in\mathbb{N}_{0}}$, and the fundamental representation $u^{(1)}$ is given by
@@ 4. Application to twisted rapid decay properties
Let us suppose that $\mathbb{G}$ is a matrix quantum group, and denote by ||$\cdot$|| ⋅ | the natural length function on $\text{Irr}(\mathbb{G})$ and by
@@
Since the twisted Fourier transforms $\mathcal{F}^{(x)}_{p},\mathcal{F}^{(x)}_{p^{\prime}}:\text{Pol}(\mathbb{G})\rightarrow c_{00}(\widehat{\mathbb{G}})$ are bijective, we have
@@
Let us suppose that $\mathbb{G}$ is a compact matrix quantum group and ||$\cdot$|| ⋅ | is the natural length function on $\text{Irr}(\mathbb{G})$. If there exists a polynomial $P$ such that
@@ 5. Characterization of boundedness of $\mathcal{F}^{(x)}_{p}$
Recall that Theorem 3.6 implies $[0,1]\subseteq I(\mathbb{G},p)$ for all $21\leq p<$ 2. This Section is devoted to characterizing $I(\mathbb{G},p)$ for non-Kac compact quantum groups $\mathbb{G}$. Subsection 5.1 focuses on non-Kac and coamenable cases, and Subsection 5.2 focuses on non-Kac and non-coamenable free orthogonal quantum groups $O_{F}^{+}$, respectively.
@@ 5.1. The case of non-Kac and coamenable compact quantum groups
The main result of this subsection is that $[0,1]=I(\mathbb{G},p)$ holds for all $21\leq p<$ 2 under the assumption that $\mathbb{G}$ is non-Kac with a polynomial growth on $\widehat{\mathbb{G}}$. To prove this, we begin with the following lemma.
@@
Let $21\leq p<$ 2, $x\in\mathbb{R}$, and $\alpha\in\text{Irr}(\mathbb{G})$. Suppose that
@@
Let $\mathbb{G}$ be a general compact quantum group, $21\leq p<$ 2 and $x\in\mathbb{R}$. Then (1) $\Rightarrow$ (2) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (5) holds in the following list.
@@
which is true only when $-x\leq 0$ and $x-1\leq 0$, i.e. $x\in[0,1]$. ∎
@@
Let $G$ be a simply connected semisimple compact Lie group and $10<q<$ 1. Then $I(G_{q},p)=[0,1]$ holds for any $21\leq p<$ 2.

### v2

@@ Abstract.
We introduce an analytic family of twisted Fourier transforms $\left\{\mathcal{F}^{(x)}_{p}\right\}_{x\in\mathbb{R},p\in[1,2)}$ for non-Kac compact quantum groups and establish a sharpened form of the Hausdorff-Young inequality in the range $0\leq x\leq 1$. Furthermore, we prove that the range $0\leq x\leq 1$ is both necessary and sufficient for the boundedness of $\mathcal{F}^{(x)}_{p}$ under the assumption of sub-exponential growth on the dual discrete quantum group. We also show that the range of boundedness of $\mathcal{F}^{(x)}_{p}$ can be strictly extended beyond $[0,1]$ for certain non-Kac and non-coamenable free orthogonal quantum groups. As applications, we derive a stronger form of the twisted rapid decay property for polynomially growing non-Kac discrete quantum groups, including the duals of the Drinfeld-Jimbo $q$-deformations, and construct an explicit contractive, but non-completely bounded, representation of the convolution algebra of any non-Kac free orthogonal quantum group.
@@ 1. Introduction
[…]
@@
[…]
@@
$I(\mathbb{G},p)$ $=\left\{x\in\mathbb{R}~\Bigr|~\mathcal{F}^{(x)}_{p}:L^{p}(\mathbb{G})\rightarrow\ell^{p^{\prime}}(\widehat{\mathbb{G}})\text{ is bounded}\right\}.$ (1.4)
Note that (1.2) guarantees that $0,1]\subseteq I(\mathbb{G},p)$ for any compact quantum group $\mathbb{G}$ and $1\leq p<2$. Subsection 5.1 focuses on non-Kac and coamenable compact quantum groups, and shows that equality $0,1]=I(\mathbb{G},p)$ holds for all $1\leq p<2$ under the assumption of sub-exponential growth on $\widehat{\mathbb{G}}$ (Theorem 5.2). This class includes the Drinfeld-Jimbo $q$-deformations $G_{q}$. Subsection 5.2 focuses on non-Kac and non-coamenable free orthogonal quantum groups $O_{F}^{+}$. Using the recently established Haagerup inequality on $O_{F}^{+}$ [BVY21, You22], we show that the inclusion $0,1]\subsetneq I(O_{F}^{+},p)$ is strict for certain non-Kac and non-coamenable free orthogonal quantum groups $O_{F}^{+}$ satisfying a technical assumption $\left\|F\right\|_{op}^{6}<\frac{d_{1}+\sqrt{d_{1}^{2}-4}}{2}$ where $d_{1}=\text{Tr}(F^{*}F)$ (Corollary 5.7). A summary of the main conclusions regarding $I(\mathbb{G},p)$ is presented in the following table:
@@
$1\leq p<2$ $I(\mathbb{G},p)=\mathbb{R}$ $I(G_{q},p)=[0,1]$ $I(O_{F}^{+},p)\supsetneq[0,1]$
@@
[…]
@@ 2.1. Compact quantum groups
[KS18, Section 6] Let $\mathbb{G}$ be a compact matrix quantum group of non-Kac type and let $|\cdot|$ be the natural length function on $\text{Irr}(\mathbb{G})$. Then there exists a sequence $(\alpha_{k})_{k\in\mathbb{N}}\subseteq\text{Irr}(\mathbb{G})$ such that
@@ 2.2. Noncommutative $L^{p}$-spaces
$\iota(a)](x)=h(xa),~x\in L^{\infty}(\mathbb{G}).$ (2.14)
@@
Let us denote by $S=\left\{z\in\mathbb{C}:0<\text{Re}(z)<1\right\}$ and by $\mathcal{C}(A)$ the set of all continuous bounded functions $f:\overline{S}\rightarrow M_{n}(\mathbb{C})$ such that $f|_{S}:S\rightarrow M_{n}(\mathbb{C})$ is analytic and $f(\frac{1}{p})=A$. The finite dimensional Schatten $p$-space $S^{p}_{n}$ is the complex interpolation space $(M_{n},S^{1}_{n})_{\theta}$ with $\theta=\frac{1}{p}$ in the sense that
@@ 2.3. Discrete quantum group and Fourier transform
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3. Twisted Fourier transforms
[…]
@@
[…]
@@
[…]
@@ 3.1. Unboundedness for the cases $p>2$
[…]
@@
[…]
@@
In this case, the numerator is given by $(Q_{\alpha})_{jj}(Q_{\alpha})_{ii}^{-1}]^{(\frac{1}{2}-\frac{1}{p})x}\cdot\left(\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right)^{\frac{1}{p}}$ by Lemma 3.2, and we have
@@
3.2. Boundedness for the cases $1\leq p<2$ and $0\leq x\leq 1$
In this Subsection, we will focus on the cases $1\leq p<2$. If $\mathbb{G}$ is of Kac type, then we have $\mathcal{F}^{(x)}_{p}=\mathcal{F}^{(0)}_{p}$ for all $x\in\mathbb{R}$, so we obtain $\mathbb{R}=I(\mathbb{G},p)$ for all $1\leq p<2$ thanks to the standard Hausdorff-Young inequality.
@@ 3.2. Boundedness for the cases $1\leq p<2$ and $0\leq x\leq 1$
The main result of this Section is that $0,1]\subseteq I(\mathbb{G},p)$ for all $1\leq p<2$. Let us begin with the case $p=1$ in the following theorem.
@@
In particular, we have $0,1]\subseteq I(\mathbb{G},1)$.
@@
[…]
@@
[…]
@@
Let $1\leq p<2$. For each $\alpha\in\text{Irr}(\mathbb{G})$, let us take indices $i=i(\alpha),j=j(\alpha)$ such that $(Q_{\alpha})_{ii}\geq(Q_{\alpha})_{jj}$. Then, for any $f\sim\sum_{\alpha\in\text{Irr}(\mathbb{G})}a_{\alpha}u^{\alpha}_{ij}\in L^{p}(\mathbb{G})$ and $x\leq y$, we have
@@
$=\left\{\begin{array}[]{lll}\left(\sum_{\alpha\in\text{Irr}(\mathbb{G})}|a_{\alpha}|^{p^{\prime}}\left[\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right]^{p^{\prime}-1}\left[(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}\right]^{\frac{p^{\prime}}{2}-1}\right)^{\frac{1}{p^{\prime}}}&,~1<p<2\\ \sup_{\alpha\in\text{Irr}(\mathbb{G})}\frac{|a_{\alpha}|}{d_{\alpha}}[(Q_{\alpha})_{jj}(Q_{\alpha})_{ii}]^{-\frac{1}{2}}&,~p=1.\end{array}\right$. (3.54)
@@
$=\left\{\begin{array}[]{ll}\left(\sum_{\alpha\in\text{Irr}(\mathbb{G})}\left[(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}\right]^{(\frac{p^{\prime}}{2}-1)x}\left[\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}\right]^{p^{\prime}-1}\left|a_{\alpha}\right|^{p^{\prime}}\right)^{\frac{1}{p^{\prime}}}&,~1<p<2\\ \sup_{\alpha\in\text{Irr}(\mathbb{G})}\left[(Q_{\alpha})_{jj}^{-1}(Q_{\alpha})_{ii}\right]^{\frac{x}{2}}\cdot\frac{(Q_{\alpha})_{ii}^{-1}}{d_{\alpha}}|a_{\alpha}|&p=1.\end{array}\right$. (3.58)
@@
Let $\mathbb{G}=SU_{q}(2)$ with $0<q<1$. Then $\text{Irr}(SU_{q}(2))$ is identified with $\left\{u^{(k)}\right\}_{k\in\mathbb{N}_{0}}$, and the fundamental representation $u^{(1)}$ is given by
@@ 4. Application to twisted rapid decay properties
Let us suppose that $\mathbb{G}$ is a matrix quantum group, and denote by $|\cdot|$ the natural length function on $\text{Irr}(\mathbb{G})$ and by
@@
[…]
@@
Let us suppose that $\mathbb{G}$ is a compact matrix quantum group and $|\cdot|$ is the natural length function on $\text{Irr}(\mathbb{G})$. If there exists a polynomial $P$ such that
@@ 5. Characterization of boundedness of $\mathcal{F}^{(x)}_{p}$
Recall that Theorem 3.6 implies $0,1]\subseteq I(\mathbb{G},p)$ for all $1\leq p<2$. This Section is devoted to characterizing $I(\mathbb{G},p)$ for non-Kac compact quantum groups $\mathbb{G}$. Subsection 5.1 focuses on non-Kac and coamenable cases, and Subsection 5.2 focuses on non-Kac and non-coamenable free orthogonal quantum groups $O_{F}^{+}$, respectively.
@@ 5.1. The case of non-Kac and coamenable compact quantum groups
The main result of this subsection is that $0,1]=I(\mathbb{G},p)$ holds for all $1\leq p<2$ under the assumption that $\mathbb{G}$ is non-Kac with a polynomial growth on $\widehat{\mathbb{G}}$. To prove this, we begin with the following lemma.
@@
Let $1\leq p<2$, $x\in\mathbb{R}$, and $\alpha\in\text{Irr}(\mathbb{G})$. Suppose that
@@
Let $\mathbb{G}$ be a general compact quantum group, $1\leq p<2$ and $x\in\mathbb{R}$. Then (1) $\Rightarrow$ (2) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (5) holds in the following list.
@@
which is true only when $x\leq 0$ and $x-1\leq 0$, i.e. $x\in[0,1]$. ∎
@@
Let $G$ be a simply connected semisimple compact Lie group and $0<q<1$. Then $I(G_{q},p)=[0,1]$ holds for any $1\leq p<2$.